Arthur can do a job in 30 min., Bonnie can do it in 40 min., Claire can do it in 60 min. How long will it take them if they work together?
Answer by richwmiller(17219)

It has to be less than 30 minutes since Arthur can do it alone in 30 minutes.
and Bonnie in 40 minutes so they should be able to do it in abt 17 minutes.
2x+1.5x+1x=1 hour
4.5x=60
x=60/4.5=600/45=13 minutes 20 seconds